A civil engineer takes a public ride going to his firm. The average time for this trip is 24 minutes, with a standard deviation of 3.80 minutes. Find the probability that 2 of the next 3 trips will take at least 1/2 hour.


1
Expert's answer
2021-05-24T05:31:02-0400

let's find the probability that a trip will take at least 1/2 hour

"P(X\ge30)=1-P
(X<30)"


"=1-P(\frac{X-\mu}{\sigma}<\frac{30-24}{3.8})"

"=1-P(Z<1.58)"


"=1-0.9429"


"=0.0571"